OpenStudy (anonymous):

A rectangular kitchen table has an area of 3x2 + 32x + 45. Which of the following could be the length of the table? 3x + 5 3x + 9 x + 5 There is not enough information given.

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ 3x^2 + 32x + 45= (3x+5)(x+9) }\)

OpenStudy (anonymous):

I don't understand..

OpenStudy (solomonzelman):

`Area = w × ℓ`

OpenStudy (solomonzelman):

This is the same thing, see ?

OpenStudy (anonymous):

v.v... i give up..

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ 3x^2 + 32x + 45 =\color{red}{(3x+5) }\times \color{green}{(x+9) } }\) \(\normalsize\color{blue}{ 3x^2 + 32x + 45 =\color{red}{w } \times \color{green}{l } }\)

OpenStudy (solomonzelman):

3x²+32+45 is the area.

OpenStudy (anonymous):

so... D?

OpenStudy (solomonzelman):

omg bbbye

OpenStudy (anonymous):

......